OpenStudy (anonymous):

For questions 1-2, convert to logarithmic equations. For example, the logarithmic form of "23 = 8" is "log2 8 = 3". 1. 9 3/2 = 27 2. ex = 7 For questions 3-4, write the logarithmic equation in exponential form. For example, the exponential form of "log5 25 = 2" is "52 = 25". 3. log3 81 = 4 4. log6 216 = 3

OpenStudy (anonymous):

1.log9 27=3/2; 2. ln7=x 3. 3^4=81, 4. 6^3=216

OpenStudy (anonymous):

1. log9 27 = log (9*9^(1/2)) = log9 9 + log9 9^(1/2) = 1 + 1/2 = 3/2

OpenStudy (anonymous):

e^x = 7 2. x= ln 7 = 1.9459

OpenStudy (anonymous):

3. log3 81 = log3 3^4 = 4 so 3^4= 81

OpenStudy (anonymous):

4. log6 216 = log6 6^3 = 3 so 6^3 = 216